Assume the given amount of RbF fully dissociates in water. Calculate the molarity of a solution if 138.42 g of RbF are dissolved into a solution of water that has a final volume of 2.00 L.


1
Expert's answer
2021-04-29T07:18:00-0400

Molar mass of RbF = 104.47 g/mol

Mass of RbF = 138.42g

number of moles = 138.42/104.47 = 1.32 moles


Volume = 2.00 L


molarity = number of moles/ Volume = "\dfrac{1.32}{2.00}" = 0.66 mol/L = 0.66 M